OpenStudy (anonymous):

What is the value of expression? 5−8 • 5−4 A. 532 B. 54 C. 5−4 D. 5−12

OpenStudy (alexandervonhumboldt2):

first thing you do is multiplication 8*5. 5-40-4=now subtract=-35-4=???

OpenStudy (anonymous):

31

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

c

Nnesha (nnesha):

please use `^` for exponents so your question is \[\huge\rm 5^{-8} \times 5^{-4}\] now apply the exponent rule when we multiply same bases we should `add` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

= 5 12

Nnesha (nnesha):

-8-4 = ??

Nnesha (nnesha):

both numbers are negative....

OpenStudy (anonymous):

so 5^-12

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

thx